RC meeting and Run Plan, Friday November 26th
From SBS wiki
RC Meeting Information
RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934
Accelerator and Hall Status
- We are cycling through HCal Calib run and production runs, all detector-study run were finished.
- Beam is stable until the end of Thursday swing shift, roughly around 11:30 PM MCC had trouble to deliver 15 uA beam (10 uA Max).
- Then the Hall A Moller Magnet failed at 3 AM during Friday OWL shift. See a MCC Log for this issue: https://logbooks.jlab.org/entry/3950911, Optic experts developed a new beam optics plan to bypass this Moller Magnet, they said they've done this before.
- During the Friday Day shift, we mainly were establishing beam into Hall, David Flay is on remote support, see a log entry by David for this work: https://logbooks.jlab.org/entry/3951014, As of now, we have back to production run.
- There's one issue that needs to note is our trigger rate, there are a couple early production runs where we set BBCal threshold to 400mV, this corresponds to 2 GeV energy cut in BBCal. But with this threshold we have around 8kHz trigger rate, and 42% live time. Data for these runs might have some issue. We increased the threshold to 450mV to give DAQ ~90% live time.
- There's a trigger configuration change, Brian and Alex suggested we add HCal LED trigger, so starting from run 12318 (Swing shift) all the production runs are using ps1=0, ps6=0, others = -1. Previously only ps1=0.
Pending and New issues
- GEM layer J0 common mode is drifting higher again, no progress on this
- HODO scope HV issue: expert confirmed that this is not a hardware issue, see log: https://logbooks.jlab.org/entry/3950539 we are fine for experiment, but this may need to be fixed later.
- Do we need to check BBCal calibration regarding the issue from trigger rate (8K at 400 mV, 2GeV)?
- J0 Bottom chamber has APVs unstable, causing dead area in the chamber and also DAQ hung. need fix. These APVs are on INFN J0 fiber 17, 19. As of now, it seems this should not be a reason to stop beam and debug.
Resolved Issues
- The troubling Moller Quad is bypassed. Hopefully we won't have any beam trip from now on.
Subsystem Updates
- HCAL:
- BBCal:
- Timing hodoscope:
- Grinch:
- GEMS:
- Beamline:
- DAQ:
- Replay analysis:
Run Plan
We will continue the production SBS-11 runplan as listed in this link: https://sbs.jlab.org/wiki/index.php/Main_Page